# The finite groups with three automorphism orbits

### v1

@@ 4.1. ${\mathrm{SL}}(3,q)\lhd G$
Recall that the automorphism $\xi\in{\mathrm{Aut}}(N)$, given in Proposition 2.5 (ii), has order $p^{n}-1$ and $\langle\xi\rangle$ acts faithfully on $V$. Hence, $\langle\xi\rangle$ induces a subgroup $H\leqslant X$ of order $p^{n}-1$ and $W^{H}=W$ by Lemma 2.10 (ii). Note that there exists $G\leqslant X$ such that $H<G$ and $G\cong{\mathrm{GL}}(3,q)$ with $q=p^{n/3}$. By Lemma 2.10 (i), 3.2 and 3.5, ${\mathbf{Z}}({\mathcal{H}})=A\oplus B$ as ${\mathbb{F}}_{p}G$-modules such that
  1. (i)
    $A\cong\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ as ${\mathbb{F}}_{p}G$-modules with $U$ the natural ${\mathbb{F}}_{q}G$-module; and
  2. (ii)
    $G$ acts transitively on non-zero vectors of ${\mathbf{Z}}({\mathcal{H}})/B\cong A$.
Thus, $N_{0}:={\mathcal{H}}/B$ is a $3$-orbit $p$-group with $|N_{0}|=|{\mathbf{Z}}(N_{0})|^{2}$ by Lemma 2.8 (iii) and 2.10 (ii). We claim that $W=B$, and then $N=N_{0}$ is a $3$-orbit $p$-group with ${\mathrm{Aut}}(N)^{V}\cong{\Gamma\mathrm{L}}(3,q)$ by Lemma 4.1.
Note that $B$ is an ${\mathbb{F}}_{p}H$-submodule of ${\mathbf{Z}}({\mathcal{H}})\cong\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$. Hence, $B=W$ if and only if ${\mathbf{Z}}(N)$ and $A$ are isomorphic ${\mathbb{F}}_{p}H$-modules deduced by Lemma 4.2. Let $h\in H$ induced by $\xi$, and let $\chi_{{\mathbf{Z}}}$ and $\chi_{A}$ be characters of $h$ on ${\mathbf{Z}}(N)$ and $A$, respectively. It suffices to show that $\chi_{{\mathbf{Z}}}=\chi_{A}$. Let $\chi_{V}$ be the character of $h$ on $V$. By definition of $\xi$, we have
$\chi_{V}=\mathrm{Tr}_{{\mathbb{F}}_{p^{n}}/{\mathbb{F}}_{p}}(\lambda)\mbox{ and }\chi_{{\mathbf{Z}}}=\mathrm{Tr}_{{\mathbb{F}}_{p^{n}}/{\mathbb{F}}_{p}}(\lambda\cdot\lambda^{\theta}),$
where $\lambda$ is a generator of ${\mathbb{F}}_{p^{n}}^{\times}$. Note that $A\cong\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ as ${\mathbb{F}}_{p}H$-modules. Let $\mathcal{U}$ be the natural ${\mathbb{F}}_{p^{n}}H$-module. Then $\mathcal{U}\cong U$ as ${\mathbb{F}}_{q}H$-modules. Hence, Lemma 3.2 deduces that $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)\cong\mathcal{U}\otimes_{{\mathbb{F}}_{p^{n}}}\mathcal{U}^{\theta}$ as ${\mathbb{F}}_{q}H$-modules since $\theta$ is a generator of ${\mathrm{Gal}}({\mathbb{F}}_{p^{n}}/{\mathbb{F}}_{q})$. Then we have
$\chi_{A}=\mathrm{Tr}_{{\mathbb{F}}_{p^{n}}/{\mathbb{F}}_{p}}(\lambda\cdot\lambda^{\theta})=\chi_{{\mathbf{Z}}}.$
@@
This yields that $W=B$. Therefore, $N$ is a $3$-orbit group and ${\mathrm{Aut}}(N)^{N/{\mathbf{Z}}(N)}\cong{\Gamma\mathrm{L}}(3,q)$ with $q=p^{n/3}$. ∎
@@ 4.2. ${\mathrm{Sp}}(2m,p^{n})\lhd G$
Now we treat the case that ${\mathrm{Sp}}(2m,q)\lhd G$. The following lemma shows that these groups are factor groups of $q^{1+2m}_{+}$.
@@
Let $q:=p^{n}$ with $p$ an odd prime, and let $N$ be a special $p$-group of exponent $p$ with $N/{\mathbf{Z}}(N)\cong{\mathbb{Z}}_{p}^{2mn}$ for some integer $m$. The following statements hold.
@@
Let ${\mathcal{H}}:={\mathcal{H}}_{2mn}$, and let $V:={\mathcal{H}}/{\mathbf{Z}}({\mathcal{H}})$. Then $X:={\mathrm{Aut}}({\mathcal{H}})^{V}={\mathrm{GL}}(V)\cong{\mathrm{GL}}(2mn,p)$ and ${\mathbf{Z}}({\mathcal{H}})$ can be identify with $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ as ${\mathbb{F}}_{p}X$-modules by Lemma 2.10 (i). Note that $N={\mathcal{H}}/W$ for some $W<{\mathbf{Z}}(N)$. Then $W^{g}=W$ for $g\in G:={\mathrm{Aut}}(N)^{V}$ by Lemma 2.10 (ii). Suppose that $G_{0}\leqslant G$ isomorphic to ${\mathrm{Sp}}(2m,q)$ and $|{\mathbf{Z}}(N)|=|\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W|=q$. Note that $q^{1+2m}_{+}={\mathcal{H}}/W_{0}$ for some $W_{0}\leqslant{\mathbf{Z}}({\mathcal{H}})$ and $X_{W_{0}}$ contains a subgroup $G_{1}$ isomorphic to ${\mathrm{Sp}}(2m,q)$ by Lemma 2.10 (ii) and 2.19. Then there exists $x\in X$ such that $G_{1}^{x}=G_{0}$, and then $q^{1+2m}_{+}\cong{\mathcal{H}}/W_{0}^{x}$ with $G_{0}\leqslant X_{W_{0}^{x}}$. Thus, we may assume that $G_{1}=G_{0}$. It suffices to show that $W_{0}=W$. By Lemma 3.7 (iii), $S\subseteq W$ and $S\subseteq W_{0}$ where $S$ is an ${\mathbb{F}}_{p}G$-submodule of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ given in Lemma 3.7. Note that $\dim S=\dim\Lambda^{2}{{\mathbb{F}}_{q}}(V)-n$, it follows that $S=W=W_{0}$ as $|{\mathbf{Z}}({\mathcal{H}})/S|=|{\mathbf{Z}}(N)|=|{\mathbf{Z}}(q^{1+2m}_{+})|=p^{n}$. Thus, $N\cong q^{1+2m}_{+}$, as in part (i).
Suppose that $G_{0}\leqslant G$ isomorphic to ${\mathrm{Sp}}(2m,q)$ and $|{\mathbf{Z}}(N)|=|\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W|\leqslant q$. With the similar arguments in part (i), we obtain that $S\leqslant W$. Then $N$ is a quotient of ${\mathcal{H}}/S\cong q^{1+2m}_{+}$ as in part (ii).
Let $N:=q^{1+2m}_{+}$. If $q=p$, then ${\mathrm{Aut}}(N)\cong p^{2m}{:}\mathrm{CSp}(2m,q)$ given by Winter [31]. Suppose that $n\geqslant 2$. Then $G$ is non-solvable by Lemma 2.19. Note that $G$ acts transitively on non-zero vectors of both $V$ and $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$ and $|V|\neq|\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W|$. Then $G$ has a normal subgroup isomorphic to ${\mathrm{Sp}}(2m^{\prime},p^{n^{\prime}})$ with $2m^{\prime}n^{\prime}=2mn$ by Theorem A. Since $G$ has a subgroup isomorphic to ${\mathrm{Sp}}(2m,q)$ Lemma 2.19, we have $n^{\prime}\leqslant n$. If $n^{\prime}<n$, then $|\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W|<q$ by Theorem A, a contradiction. Thus, $G$ has a normal subgroup isomorphic to ${\mathrm{Sp}}(2m,q)$. Therefore, $G={\mathrm{C}\Gamma\mathrm{Sp}}(2m,q)$ and we obtain part (iii) by Lemma 2.8 (i) and 2.11. ∎
We remark that the $1$-subspace stabilizer $G_{\alpha}$ of $G:={\mathrm{Sp}}(2m+2,q)$ is of form
$q^{1+2m}{:}({\mathrm{Sp}}(2m,q)\times{\mathrm{GL}}(1,q)),$
and $N:=\textbf{O}_{p}(G_{\alpha})\cong q^{1+2m}$ is a special $p$-group with $|{\mathbf{Z}}(N)|=q$ and ${\mathrm{Sp}}(2m,q)\leqslant{\mathrm{Aut}}(N)^{N/{\mathbf{Z}}(N)}$ (see [30, Section 3.5.4]). These facts together with Lemma 4.4 (i) yield that $N\cong q^{1+2m}_{+}$, which is Theorem C (iii).
@@
Let $G:={\mathrm{PSp}}(2m+2,q)$ acts naturally on $1$-spaces of ${\mathbb{F}}_{q}^{2m+2}$ where $q=p^{n}$ for odd prime $p$. Then $\textbf{O}_{p}(G_{\alpha})\cong q^{1+2m}_{+}$ for any stabilizer $G_{\alpha}$.
@@ 5. Subfield hyperplanes and quotients of $q^{1+2m}_{+}$
Let $q_{0}=p^{d}$ and $q=p^{n}$ with $d\mid n$. Note that ${\mathrm{Sp}}(2m,q)\leqslant{\mathrm{Sp}}(2mn/d,q_{0})$ for any positive integer $m$. Lemma 4.4 (ii) deduces that there exists $U<{\mathbf{Z}}(q^{1+2m}_{+})$ such that
@@
Clearly that $U$ is a hyperplane of ${\mathbb{F}}_{p^{d}}^{n/d}$ as ${\mathbb{F}}_{p^{d}}$-space. Thus, $U\in\Sigma$ and $U^{x^{\ell}}=U$. In particular, we have
@@
Let $N:=q^{1+2m}_{+}$, and let $N_{0}:=(q_{0})^{1+2mn/d}_{+}$. By Lemma 4.4 (iii), we have ${\mathrm{Aut}}(N_{0})^{N_{0}/{\mathbf{Z}}(N_{0})}\cong{\mathrm{C}\Gamma\mathrm{Sp}}(2mn/d,q_{0})$ which has a subgroup isomorphic to ${\mathrm{Sp}}(2m,q)$. Hence, $N_{0}\cong N/U_{0}$ for some $U_{0}<{\mathbf{Z}}(N)$ by Lemma 4.4 (i). Set $H:={\mathrm{Aut}}(N)^{{\mathbf{Z}}(N)}$. Then $H\cong{\Gamma\mathrm{L}}(1,q)$ by Lemma 2.19. Let $x$ be the element of order $q-1$ in $H$, and let $\delta$ be a preimage of $x$ in ${\mathrm{Aut}}(N)^{N/{\mathbf{Z}}(N)}$. Then $\delta^{\ell}\in{\mathrm{Aut}}(N_{0})^{N_{0}/{\mathbf{Z}}(N_{0})}\cong{\mathrm{C}\Gamma\mathrm{Sp}}(2mn/d,q_{0})$ for $\ell=\frac{q-1}{q_{0}-1}$ (see [18, page 112]). Hence, $U_{0}$ is invariant under $\delta^{\ell}$, equivalently, $U_{0}^{x^{\ell}}=U_{0}$. Thus, $U_{0}<{\mathbf{Z}}(N)$ is a subfield hyperplane of with respect to ${\mathbb{F}}_{p^{d}}$ by definition.
@@
Let $N=q^{1+2m}_{+}$ with $q=p^{n}$, and let $U<{\mathbf{Z}}(N)$. Suppose that $U$ contains a subfield hyperplane $U_{0}$ of ${\mathbf{Z}}(N)$ with respect to ${\mathbb{F}}_{q_{0}}$ with $q_{0}=p^{d}$,. By Theorem C (i), we have
@@
where $U_{1}:=U_{0}/U$. Thus, we may focus on the case where $U$ contains no subfield hyperplanes, as in the following example.
@@
Let $V:=q^{1+2m}_{+}/{\mathbf{Z}}(q^{1+2m}_{+})$, and identify $N/{\mathbf{Z}}(N)$ with $V$. By Lemma 4.4 (iv), we have $G:={\mathrm{Aut}}(q^{1+2m}_{+})^{V}\cong{\mathrm{C}\Gamma\mathrm{Sp}}(2m,q)$. Let $G_{0}$ be the normal subgroup of $G$ isomorphic to ${\mathrm{Sp}}(2m,q)$. Then $G_{0}$ acts trivially on ${\mathbf{Z}}(q^{1+2m}_{+})$ by Lemma 2.19. Thus, $G_{0}$ is a subgroup of ${\mathrm{Aut}}(N)^{V}$ as $U^{G_{0}}=U$. Hence, ${\mathrm{Aut}}(N)^{V}$ is transitive on non-zero vectors of $N/{\mathbf{Z}}(N)$.
Let $H:={\mathrm{Aut}}(q^{1+2m}_{+})^{{\mathbf{Z}}(q^{1+2m}_{+})}\cong G/G_{0}\cong{\Gamma\mathrm{L}}(1,q)$. Then $H_{U}\leqslant{\mathrm{Aut}}(N)^{{\mathbf{Z}}(N)}$. Since $H_{U}$ is transitive on non-zero vectors of ${\mathbf{Z}}(N)$, we have ${\mathrm{Aut}}(N)^{{\mathbf{Z}}(N)}$ is transitive on non-identity elements of ${\mathbf{Z}}(N)$. By Lemma 2.8 (iv), we obtain that $N$ is a $3$-orbit group.
@@
Let $N$ be a finite $3$-orbit group. Note that the set of orders of elements of $N$ can only be $\{1,p,q\}$, $\{1,p,p^{2}\}$ and $\{1,p\}$ for some primes $p$ and $q$. Hence, $N$ is either a $\{p,q\}$-group or a $p$-group. If $N$ is a $\{p,q\}$-group, then Lemma 2.2 shows that $N$ is in the first row of Table 1. If $N$ is an abelian $p$-group, then $N$ is in the second row by Lemma 2.3. If $N$ is a non-abelian $2$-group, then $N$ is one of the groups in rows 3-5 by Theorem 2.6.
Assume that $N$ is a non-abelian $p$-group for odd prime $p$. Then $N$ is a special $p$-group of exponent $p$ by Lemma 2.7. If ${\mathrm{Aut}}(N)$ is solvable, then $N\cong q^{1+2}_{+}$ by Theorem 2.16 and is in the 7-th row of Table 1. If ${\mathrm{Aut}}(N)$ is non-solvable, then $G:={\mathrm{Aut}}(N)^{N/{\mathbf{Z}}(N)}$ is non-solvable by Lemma 2.8 (i). Lemma 2.8 (iii) shows that $G$ acts transitively on non-identity elements of both $N/{\mathbf{Z}}(N)$ and ${\mathbf{Z}}(N)$, and Lemma 2.8 (iv) shows that ${\mathbf{Z}}(N)$ is isomorphic to a quotient ${\mathbb{F}}_{p}G$-module of $\Lambda_{{\mathbb{F}}_{p}}^{2}(N/{\mathbf{Z}}(N))$. By Theorem A, one of the following cases hold.
  1. (1)

### v2

@@ 4.1. ${\mathrm{SL}}(3,q)\lhd G$
Note that $N$ is a special $p$-group of exponent $p$. We assume that $N={\mathcal{H}}/W$ with ${\mathcal{H}}={\mathcal{H}}_{n,p}$ and $W<{\mathbf{Z}}({\mathcal{H}})$. Let $V={\mathcal{H}}/{\mathbf{Z}}({\mathcal{H}})=N/{\mathbf{Z}}(N)$, and let $X={\mathrm{Aut}}({\mathcal{H}})^{V}={\mathrm{GL}}(V)\cong{\mathrm{GL}}(n,p)$. Recall that the automorphism $\xi\in{\mathrm{Aut}}(N)$, given in Proposition 2.5 (ii), has order $p^{n}-1$ and $\langle\xi\rangle$ acts faithfully on $V$. Hence, $\langle\xi\rangle$ induces a subgroup $H\leqslant X$ of order $p^{n}-1$ and $W^{H}=W$ by Lemma 2.10 (ii). Then there exists $G\leqslant X$ such that $H<G$ and $G\cong{\mathrm{GL}}(3,q)$ with $q=p^{n/3}$. By Lemma 3.2, there exists a unique ${\mathbb{F}}_{p}G$-submodule $W_{1}<\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ such that $G$ acts transitively on non-zero vectors of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W_{1}$. We identify $W_{1}$ as a subgroup of ${\mathbf{Z}}({\mathcal{H}})$. Then $N_{1}={\mathcal{H}}/W_{1}$ is a $3$-orbit group as given in the proof of Lemma 4.1 (i). It suffices to show that $W_{1}=W$, and then $N=N_{1}$ is a $3$-orbit group with ${\mathrm{Aut}}(N)\cong p^{n}{:}{\Gamma\mathrm{L}}(3,p^{n/3})$ as given in Lemma 4.1 (iii).
Remark that each of $\Lambda_{{\mathbb{F}}_{p}}(V)$, $W$ and $W_{1}$ is an ${\mathbb{F}}_{p}H$-module and is a direct sum of non-isomorphic irreducible ${\mathbb{F}}_{p}H$-submodules by Lemma 4.2. Thus, we have that $W_{1}=W$ if and only if $\Lambda_{{\mathbb{F}}_{p}}(V)/W\cong\Lambda_{{\mathbb{F}}_{p}}(V)/W_{1}$ as ${\mathbb{F}}_{p}H$-modules. It is clear that both of these two ${\mathbb{F}}_{p}H$-modules are irreducible. Let $\chi$ and $\chi_{1}$ be the characters of $\Lambda_{{\mathbb{F}}_{p}}(V)/W$ and $\Lambda_{{\mathbb{F}}_{p}}(V)/W_{1}$. We only need to verify that $\chi(\xi)=\chi_{1}(\xi)$. Let $U$ be the natural ${\mathbb{F}}_{q}H$-module with $q=p^{n}$. Assume that $U\cong V$ as ${\mathbb{F}}_{p}H$-modules. Let $\lambda$ be the ${\mathbb{F}}_{q}$-character of $\xi$ on $U$. Then $\lambda$ generates ${\mathbb{F}}_{q}^{\times}$ and the character of $\xi$ acting on $V$ is $\mathrm{Tr}_{{\mathbb{F}}_{q}/{\mathbb{F}}_{p}}(\lambda)$. By definition of $\xi$, we have that
$\chi(\xi)=\mathrm{Tr}_{{\mathbb{F}}_{q}/{\mathbb{F}}_{p}}(\lambda\cdot\lambda^{\theta})=\sum_{i=1}^{n}\lambda^{p^{i}(1+p^{\frac{n}{3}})}.$
Recall that $G\cong{\mathrm{GL}}(3,p^{n/3})$. Let $\mathcal{U}$ be the natural ${\mathbb{F}}_{p^{\frac{n}{3}}}G$-module. Let $\chi_{\mathcal{U}}$ be the ${\mathbb{F}}_{\frac{n}{3}}$-character of $G$ on $\mathcal{U}$. Then
$\chi_{\mathcal{U}}(\xi)=\mathrm{Tr}_{{\mathbb{F}}_{q}/{\mathbb{F}}_{p^{n/3}}}(\lambda)=\lambda+\lambda^{p^{\frac{n}{3}}}+\lambda^{p^{\frac{2n}{3}}}.$
In the proof of Lemma 3.2, we obtain that $\Lambda_{{\mathbb{F}}_{p}}(V)/W_{1}$ is ${\mathbb{F}}_{p}G$-isomorphic to $\Lambda_{{\mathbb{F}}_{\frac{n}{3}}}^{2}(\mathcal{U})$. Thus, we have that
$\chi_{1}(\xi)$ $=\mathrm{Tr}_{{\mathbb{F}}_{p^{n/3}}/{\mathbb{F}}_{p}}\left(\frac{\chi_{\mathcal{U}}(\xi)^{2}-\chi_{\mathcal{U}}(\xi^{2})}{2}\right)=\sum_{i=1}^{\frac{n}{3}}\left(\lambda^{1+p^{\frac{n}{3}}}+\lambda^{1+p^{\frac{2n}{3}}}+\lambda^{p^{\frac{n}{3}}+p^{\frac{2n}{3}}}\right)^{p^{i}}$
$=\sum_{i=1}^{\frac{n}{3}}\left(\sum_{k=1}^{3}\left(\lambda^{p^{\frac{kn}{3}}(1+p^{\frac{n}{3}})}\right)\right)^{p^{i}}=\sum_{i=1}^{n}\lambda^{p^{i}(1+p^{\frac{n}{3}})}=\chi(\xi).$
@@
This yields that $W=W_{1}$, and therefore $N=N_{1}$ is a $3$-orbit group with ${\mathrm{Aut}}(N)^{N/{\mathbf{Z}}(N)}\cong{\Gamma\mathrm{L}}(3,p^{\frac{n}{3}})$. ∎
@@ 4.2. ${\mathrm{Sp}}(2m,p^{n})\lhd G$
Now we treat the case where ${\mathrm{Sp}}(2m,q)\lhd G$. The following lemma shows that these groups are factor groups of $q^{1+2m}_{+}$.
@@
Let $q=p^{n}$ with $p$ an odd prime, and let $N$ be a special $p$-group of exponent $p$ with $N/{\mathbf{Z}}(N)\cong{\mathbb{Z}}_{p}^{2mn}$ for some integer $m$. The following statements hold.
@@
Let ${\mathcal{H}}={\mathcal{H}}_{2mn}$, and let $V={\mathcal{H}}/{\mathbf{Z}}({\mathcal{H}})$. Then $X={\mathrm{Aut}}({\mathcal{H}})^{V}={\mathrm{GL}}(V)\cong{\mathrm{GL}}(2mn,p)$ and ${\mathbf{Z}}({\mathcal{H}})$ can be identified with $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ as ${\mathbb{F}}_{p}X$-modules by Lemma 2.10 (i). Note that $N={\mathcal{H}}/W$ for some $W<{\mathbf{Z}}(N)$. Let $G={\mathrm{Aut}}(N)^{V}$. Then $W^{g}=W$ for any $g\in G$.
Suppose that $G_{0}\leqslant G$ isomorphic to ${\mathrm{Sp}}(2m,q)$ and $|{\mathbf{Z}}(N)|=|\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W|=q$. Note that $q^{1+2m}_{+}={\mathcal{H}}/W_{0}$ for some $W\leqslant{\mathbf{Z}}({\mathcal{H}})$ and $X_{W_{0}}$ contains a subgroup $G_{1}$ isomorphic to ${\mathrm{Sp}}(2m,q)$ by Lemma 2.10 (ii) and 2.19. Then there exists $x\in X$ such that $G_{1}^{x}=G_{0}$, and then $q^{1+2m}_{+}\cong{\mathcal{H}}/W^{x}$ with $G_{0}\leqslant X_{W^{x}}$. Thus, we may assume that $W_{0}^{x}=W_{0}$ and $G_{1}=G_{0}$. It suffices to show that $W_{0}=W$. By Lemma 3.3 (iii), $S\subseteq W$ and $S\subseteq W_{0}$, where $S$ is unique determined by $G_{0}=G^{(\infty)}$. Note that $\dim S=\dim\Lambda^{2}{{\mathbb{F}}_{q}}(V)-n$, it follows that $S=W=W_{0}$ as $|{\mathbf{Z}}({\mathcal{H}})/S|=|{\mathbf{Z}}(N)|=|{\mathbf{Z}}(q^{1+2m}_{+})|=p^{n}$. Thus, $N\cong q^{1+2m}_{+}$, as in part (i).
Part (ii) can be proved with similar arguments as above by Lemma 3.3 (i).
Let $N=q^{1+2m}_{+}$. If $q=p$, then ${\mathrm{Aut}}(N)\cong p^{2m}{:}\mathrm{CSp}(2m,q)$ given by Winter [37]. Suppose that $n\geqslant 2$. Then $G={\mathrm{Aut}}(N)^{V}$ contains a subgroup isomorphic to ${\mathrm{C}\Gamma\mathrm{Sp}}(2m,q)$ by Lemma 2.19. Hence, $G$ is non-solvable. Remark that $G$ acts transitively on non-zero vectors of both $V$ and $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$ and $|V|\neq|\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W|$. Then $G$ has a normal subgroup isomorphic to ${\mathrm{Sp}}(2m^{\prime},p^{n^{\prime}})$ with $2m^{\prime}n^{\prime}=2mn$ by Theorem A. As ${\mathrm{C}\Gamma\mathrm{Sp}}(2m,q)\lesssim G$, we have that $n^{\prime}\leqslant n$. Suppose that $n^{\prime}<n$. Then $|\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W|<q$ by Theorem A, a contradiction. Thus, $G$ has a normal subgroup isomorphic to ${\mathrm{Sp}}(2m,q)$. Therefore, $G\cong{\mathrm{C}\Gamma\mathrm{Sp}}(2m,q)$ and we obtain part (iii) by Lemmas 2.8 (i) and 2.11. ∎
We remark that the $1$-subspace stabilizer $G_{\alpha}$ of $G={\mathrm{Sp}}(2m+2,q)$ has form
$q^{1+2m}{:}({\mathrm{Sp}}(2m,q)\times{\mathrm{GL}}(1,q)).$
Let $N=\textbf{O}_{p}(G_{\alpha})\cong q^{1+2m}$. Then $N$ is a special $p$-group with $|{\mathbf{Z}}(N)|=q$ and ${\mathrm{Sp}}(2m,q)\lesssim{\mathrm{Aut}}(N)^{N/{\mathbf{Z}}(N)}$ (see [36, Section 3.5.4]). These facts together with Lemma 4.4 (i) yield that $N\cong q^{1+2m}_{+}$, which is Theorem C (iii).
@@
Let $G={\mathrm{PSp}}(2m+2,q)$ where $q=p^{n}$ with odd prime $p$. Then $G$ naturally acts on $1$-spaces of ${\mathbb{F}}_{q}^{2m+2}$ and $\textbf{O}_{p}(G_{\alpha})\cong q^{1+2m}_{+}$ for any stabilizer $G_{\alpha}$.
@@ 5. Subfield hyperplanes and quotients of $q^{1+2m}_{+}$
Let $q_{0}=p^{d}$ and $q=p^{n}$ with $d\mid n$. Note that ${\mathrm{Sp}}(2m,q)\leqslant{\mathrm{Sp}}(2mn/d,q_{0})$ for any positive integer $m$. Lemma 4.4 (ii) implies that there exists $U<{\mathbf{Z}}(q^{1+2m}_{+})$ such that
@@
Clearly, $U$ is a hyperplane of ${\mathbb{F}}_{p^{d}}^{n/d}$ as ${\mathbb{F}}_{p^{d}}$-space. Thus, $U\in\Sigma$ and $U^{x^{\ell}}=U$. In particular, we have
@@
Let $N=q^{1+2m}_{+}$, and let $N_{0}=(q_{0})^{1+2mn/d}_{+}$. By Lemma 4.4 (iii), we have ${\mathrm{Aut}}(N_{0})^{N_{0}/{\mathbf{Z}}(N_{0})}\cong{\mathrm{C}\Gamma\mathrm{Sp}}(2mn/d,q_{0})$ which has a subgroup isomorphic to ${\mathrm{Sp}}(2m,q)$. Hence, $N_{0}\cong N/U_{0}$ for some $U_{0}<{\mathbf{Z}}(N)$ by Lemma 4.4 (i). Set $H={\mathrm{Aut}}(N)^{{\mathbf{Z}}(N)}$. Then $H\cong{\Gamma\mathrm{L}}(1,q)$ by Lemma 2.19. Let $x$ be an element of order $q-1$ in $H$, and let $\delta$ be a preimage of $x$ in ${\mathrm{Aut}}(N)^{N/{\mathbf{Z}}(N)}$. Then $\delta^{\ell}\in{\mathrm{Aut}}(N_{0})^{N_{0}/{\mathbf{Z}}(N_{0})}\cong{\mathrm{C}\Gamma\mathrm{Sp}}(2mn/d,q_{0})$ for $\ell=\frac{q-1}{q_{0}-1}$ (see [21, page 112]). Hence, $U_{0}$ is invariant under $\delta^{\ell}$, equivalently, $U_{0}^{x^{\ell}}=U_{0}$. Thus, $U_{0}<{\mathbf{Z}}(N)$ is a subfield hyperplane of with respect to ${\mathbb{F}}_{p^{d}}$ by definition.
@@
Let $N=q^{1+2m}_{+}$ with $q=p^{n}$, and let $U<{\mathbf{Z}}(N)$. Suppose that $U$ contains a subfield hyperplane $U_{0}$ of ${\mathbf{Z}}(N)$ with respect to ${\mathbb{F}}_{q_{0}}$ with $q_{0}=p^{d}$. By Theorem C (i), we have
@@
where $U_{1}=U_{0}/U$. Thus, we may focus on the case where $U$ contains no subfield hyperplanes, as in the following example.
@@
Let $V=q^{1+2m}_{+}/{\mathbf{Z}}(q^{1+2m}_{+})$, and identify $N/{\mathbf{Z}}(N)$ with $V$. By Lemma 4.4 (iv), we have $G={\mathrm{Aut}}(q^{1+2m}_{+})^{V}\cong{\mathrm{C}\Gamma\mathrm{Sp}}(2m,q)$. Let $G_{0}$ be the normal subgroup of $G$ isomorphic to ${\mathrm{Sp}}(2m,q)$. Then $G_{0}$ acts trivially on ${\mathbf{Z}}(q^{1+2m}_{+})$ by Lemma 2.19. Thus, $G_{0}$ is a subgroup of ${\mathrm{Aut}}(N)^{V}$ as $U^{G_{0}}=U$. Hence, ${\mathrm{Aut}}(N)^{V}$ is transitive on non-zero vectors of $N/{\mathbf{Z}}(N)$.
Let $H={\mathrm{Aut}}(q^{1+2m}_{+})^{{\mathbf{Z}}(q^{1+2m}_{+})}\cong G/G_{0}\cong{\Gamma\mathrm{L}}(1,q)$. Then $H_{U}\leqslant{\mathrm{Aut}}(N)^{{\mathbf{Z}}(N)}$. Since $H_{U}$ is transitive on non-zero vectors of ${\mathbf{Z}}(N)$, we have ${\mathrm{Aut}}(N)^{{\mathbf{Z}}(N)}$ is transitive on non-identity elements of ${\mathbf{Z}}(N)$. By Lemma 2.8 (iv), we obtain that $N$ is a $3$-orbit group.
@@
Let $N$ be a finite $3$-orbit group. Note that the set of orders of elements of $N$ can only be $\{1,p,q\}$, $\{1,p,p^{2}\}$ and $\{1,p\}$ for some primes $p$ and $q$. Hence, $N$ is either a $\{p,q\}$-group or a $p$-group.
If $N$ is a $\{p,q\}$-group, then Lemma 2.2 shows that $N$ is in the first row of Table 1. If $N$ is an abelian $p$-group, then $N$ is in the second row by Lemma 2.3. If $N$ is a non-abelian $2$-group, then $N$ is one of the groups in rows 3-5 by Theorem 2.6.
Assume that $N$ is a non-abelian $p$-group for odd prime $p$. Then $N$ is a special $p$-group of exponent $p$ by Lemma 2.7. If ${\mathrm{Aut}}(N)$ is solvable, then $N\cong q^{1+2}_{+}$ by Theorem 2.16 and is in the 7-th row of Table 1. If ${\mathrm{Aut}}(N)$ is non-solvable, then $G={\mathrm{Aut}}(N)^{N/{\mathbf{Z}}(N)}$ is non-solvable by Lemma 2.8 (i). Lemma 2.8 (iii) shows that $G$ acts transitively on non-identity elements of both $N/{\mathbf{Z}}(N)$ and ${\mathbf{Z}}(N)$, and Lemma 2.8 (iv) shows that ${\mathbf{Z}}(N)$ is isomorphic to a quotient ${\mathbb{F}}_{p}G$-module of $\Lambda_{{\mathbb{F}}_{p}}^{2}(N/{\mathbf{Z}}(N))$. By Theorem A, one of the following cases holds.
  1. (a)